OpenStudy (agent47):

@hartnn stuck on part two of that problem now.. How would I find the probability density function of the normal random variable X=3A-B

OpenStudy (agent47):

Oh and I'm given that two standard normal variables A and B are independent.

hartnn (hartnn):

\( f_{U+V}(x) = \int_{-\infty}^\infty f_U(y) f_V(x - y)\,dy = \left( f_{U} * f_{V} \right) (x) \)

hartnn (hartnn):

do you have options/choices?

OpenStudy (agent47):

Nope

hartnn (hartnn):

hmm...my first thought was to just put X=3A-B in the definition of std, normal fn, but then i saw this...

OpenStudy (agent47):

trying to find something in the book, but as usual, they never assign probability problems that are similar to the examples solved in the book...

OpenStudy (agent47):

I think I have something here... So: PDF: \[f(x)\ge0 \space for \space all \space x \in \mathbb{R} \]

OpenStudy (agent47):

and \[\int_{-\infty}^{\infty}f(x)dx=1\]

hartnn (hartnn):

those r true for any pdf

OpenStudy (agent47):

We know that: \[\int_{-\infty}^{\infty}x^2*f(x)dx=E[X^2]=1\]eh nvm, this doesn't really help

hartnn (hartnn):

yeah, mean variance has not much to do here..

OpenStudy (agent47):

Ok so the standard normal variable denoted by Z is the random variable with pdf: \[f(x)=\frac{1}{\sqrt{2\pi}}*e^{-\frac{x^2}{2}}\]